Which of the following bases produces two OH- ions per formula unit of base when dissolved in water?
Ammonia
Aluminum hydroxide
Potassium hydroxide
Calcium hydroxide
1
Expert's answer
2020-05-10T09:20:43-0400

Solution:

1) Ammonia 

When NH3 is dissolved in H2O (water) it will dissociate (dissolve) into one NH4+ ion and one OH- ion.

NH3(aq) + H2O(l) = NH4+(aq) + OH-(aq)

2) Aluminum hydroxide

When Al(OH)3 is dissolved in H2O (water) it will dissociate (dissolve) into one Al3+ ion and three OH- ions. 

Al(OH)3(s) = Al3+(aq) + 3OH-(aq)

3) Potassium hydroxide 

When KOH is dissolved in H2O (water) it will dissociate (dissolve) into one K+ ion and one OH- ion.

KOH(aq) = K+(aq) + OH-(aq)

4) Calcium hydroxide

When Ca(OH)2 is dissolved in H2O (water) it will dissociate (dissolve) into one Ca2+ ion and two OH- ions.

Ca(OH)2(s) = Ca2+(aq) + 2OH-(aq)


Answer: Calcium hydroxide produces two OH- ions when dissolved in water.